# Hierarchy Pruning for Unseen Domain Discovery in Predictive Healthcare

### v2

@@ 3.1 Step 1: Hierarchy-Guided Domain Discovery
Each domain can be viewed as a latent representation $\mathbf{r}$ sampled from a meta domain distribution $p(\cdot)$, so that we can identify $\mathbf{r}$ and then factorize $p(y|\mathbf{x})$ into $\int p(y|\mathbf{x},\mathbf{r})p(\mathbf{r}|\mathbf{x})\,d\mathbf{r}$ by approximating $q(\mathbf{r})\sim p(\mathbf{r}|\mathbf{x})$ given data samples $\mathbf{x}$. Subsequently, a domain encoder $p(\mathbf{r}|\mathbf{x})$ and a label predictor $p(y|\mathbf{x},\mathbf{r})$ are needed for inference. Here we parameterize the domain encoder $p(\mathbf{r}|\mathbf{x})$ as a network $g_{\theta}(\cdot)$ with parameter $\theta$. Since the pruned output matrix $\mathbf{M}$ (see Section 3.1) maps each training sample $\mathbf{x}$ to $\mathbf{m}$ (soft-label domain IDs), we apply $g_{\theta}(\cdot)$ to $\mathbf{m}$ to estimate the domain factor $\mathbf{r}:=g_{\theta}(\mathbf{m})$. Although $\mathbf{r}$ represents a probabilistic domain variable, we implement $g_{\theta}$ as a deterministic Multi-Layer Perceptron (MLP) for the prediction task. Next, we compute invariant features using a non-parametric function $h(\cdot):(\mathbf{r},\mathbf{p})\mapsto\mathbf{h}$, which fuses $\mathbf{r}$ (domain-level representation) and $\mathbf{p}$ (patient-level representation) as input features for the label predictor $p(y|\mathbf{x},\mathbf{r})$.
@@
$\mathcal{L}_{r}[g_{\theta}(\mathbf{m}),\bar{\mathbf{p}}]:=\mathrm{MSE}(\mathbf{r},\;\mathbb{E}[\mathbf{p}|\mathbf{m}])+\frac{\|\mathbf{r}_{\mu}-\mathbf{p}_{\mu}\|_{\mathcal{F}}^{2}}{\|\mathbf{p}_{\mu}\|_{\mathcal{F}}^{2}}.$ (6)
@@
Invariant Feature Projection Learning. In equation 6, both $\mathbf{r}$ and $\mathbf{p}$ are rescaled into a shared vector space with comparable magnitudes. Hence, we can directly apply an orthogonal projection approach (as in early studies (Bousmalis et al., 2016; Shen et al., 2022; Yang et al., 2023a)) to obtain the invariant feature $\mathbf{h}$ by subtracting the parallel component of $\mathbf{p}$ in this shared vector space. We formalize this in $h(\cdot)$ as shown in equation 7:
@@ 3.3 Training and Inference
Iterative Training. To train UdonCare, we feed each data sample $\mathbf{x}$ into the hierarchy-pruning module to obtain its latent domain $\mathbf{m}$, and then perform two cross-reference steps under a mutual learning architecture. Rather than updating the model continuously in each epoch, we adopt an iterative training strategy, which prior studies (Cui et al., 2019; Sofiiuk et al., 2022) have shown can reduce training time while maintaining comparable predictive performance. [^1] For example, we set iterations $I=3$ and epochs $N=100$ by first obtaining pretrained parameters for 40 epochs and then updating $\mathbf{M}$ iteratively every 20 epochs, yielding a total of 100 epochs. We iteratively update the model weights and regenerate domain assignments every 20 epochs in our experiment. Before each iteration, we reinitialize the parameters in $g_{\theta}(\cdot)$, because the input shape of $\mathbf{m}$ may change due to updated code-level embeddings. We also provide the pseudo-code of UdonCare in Appendix B.
Mutual Inference. After the orthogonal projection, we apply the network $q_{\xi}(\cdot)$ (operates on space $\mathbf{p}$) as a post-step to parameterize the label predictor $p(y|\mathbf{x},\mathbf{r})$; It is also available to parameterize $p(y|\mathbf{x})$ through the regular decoder network $d_{\eta}(\cdot)$ fed by patient embeddings from backbone $f_{\phi}(\cdot)$.
$p(y|\mathbf{x})\sim\hat{y}_{p}$ $=d_{\eta}(\mathbf{p})=d_{\eta}(f_{\phi}(\mathbf{x}))$
$p(y|\mathbf{x},\mathbf{r})\sim\hat{y}_{h}$ $=q_{\xi}(\mathbf{h})=q_{\xi}(h(g_{\theta}(\mathbf{m}),\mathbf{p}))$ (8)
@@ 4.1 Experiment Setups
Table 1: Statistics of MIMIC-III and MIMIC-IV datasets.
MIMIC-III 6,497 42 2.66 13.06 4.54 33.71
MIMIC-IV 49,558 55 3.66 13.38 4.70 43.89
Predictive Tasks. We evaluate our approach on four representative tasks: (1) Mortality Prediction, which determines whether a patient will pass away by a specified time horizon after discharge. This is a binary classification task. (2) Readmission Prediction, which checks if a patient will be readmitted within a predefined window (e.g., next 15 days) following discharge. This is also framed as a binary classification. (3) Diagnosis Prediction, which forecasts the set of diagnoses (ICD-9-CM codes) for the patient’s next hospital visit based on prior visits. This requires multi-label classification. (4) Drug Recommendation, which suggests a set of medications (ATC-4 codes (Nahler & Nahler, 2009)) for the upcoming visit, also formulated as multi-label classification. These tasks reflect diverse clinical needs and provide a rigorous benchmark for evaluating DG methods.
Datasets & Data Split. We conduct experiments on two publicly available EHR databases, MIMIC-III and MIMIC-IV, which are widely used in clinical prediction (Johnson et al., 2016; 2023). MIMIC-III covers ICU admissions from 2001 to 2012, while MIMIC-IV spans 2008 to 2019. To avoid overlapping time ranges with MIMIC-III, we only retain patients from the years 2013201320192019 in MIMIC-IV. For each set of experiments, we extract 6,4976{,}497 and 49,55849{,}558 patients with multiple visits ($T\geq 2$) from both datasets as shown in Table 1. Different from random data splitting, we evaluate our model’s performance across temporal gaps, following the approach in SLDG (Wu et al., 2023). We define a temporal grid based on the year of each patient’s most recent visit. Specifically, patients in MIMIC-IV (MIMIC-III) whose last visit occurred after 2017 (2010) are assigned to the target test set, while those with earlier visits are used as the source training/validation set. We also divide the dataset into training, validation, and test subsets using a fixed ratio of 75%:10%:15%.
Evaluation Metrics. Both readmission and mortality prediction are binary classification tasks, we calculate the Area Under the Precision-Recall Curve (AUPRC) and the Area Under the Receiver Operating Characteristic Curve (AUROC) scores due to the imbalanced label distribution. For the drug recommendation task, we evaluate predictions of all DG approaches by AUPRC and F1-score, following the same setting as ManyDG (Yang et al., 2023a) (i.e. $d<120$). For the diagnosis prediction tasks, we decide accurate prediction by weighted $\text{F}_{1}$ score as in Timeline and top-10 recall as in DoctorAI (Choi et al., 2016), since the former one measures the overall prediction on all classes (i.e. $d>4500$) and the latter one have concentration on positive code with low frequency.
Baselines. We compare UdonCare with two naïve baselines, five general DG baselines and two most recent clinical DG baselines: (1) Naïve Baselines: Oracle, trained backbone encoder directly on the target data, and Base, trained solely on the source data. The difference between metrics from Oracle (upper bound) and Base (lower bound) can show the distribution gaps between source and target data. (2) Typical DG Baselines: DANN (Ganin et al., 2016), CondAdv (Isola et al., 2017), MLDG (Li et al., 2018a), IRM (Arjovsky et al., 2019), and PCL (Yao et al., 2022). (3) Clinical DG Baselines: ManyDG (Yang et al., 2023a), and SLDG (Wu et al., 2023). Since drug recommendation and diagnosis prediction are multi-label classification, we drop PCL and SLDG in these two tasks due to their setting limitation. More details of the baselines can be found in Appendix C, and UdonCare is implemented as described in Appendix D.
@@
Oracle 16.73 (0.51) 70.35 (0.55) 8.46 (0.53) 68.92 (0.47) 73.42 (0.47) 69.74 (0.51) 67.37 (0.12) 66.89 (0.11)
Base 11.31 (0.62) 55.21 (0.95) 3.97 (0.47) 59.13 (0.76) 50.13 (0.88) 45.27 (0.71) 48.34 (0.24) 45.72 (0.31)
DANN 12.54 (0.55) 63.08 (0.72) 4.41 (0.45) 63.82 (0.46) 58.29 (0.63) 49.22 (0.78) 53.13 (0.11) 47.91 (0.23)
CondAdv 13.75 (0.49) 65.53 (0.57) 5.65 (0.62) 64.27 (0.68) 61.31 (0.45) 52.45 (0.51) 56.95 (0.14) 50.77 (0.19)
MLDG 13.12 (0.47) 64.48 (0.61) 4.75 (0.38) 62.75 (0.57) 60.12 (0.53) 51.03 (0.62) 55.62 (0.23) 49.57 (0.28)
IRM 13.74 (0.45) 65.21 (0.58) 4.14 (0.52) 62.36 (0.61) 60.97 (0.47) 52.02 (0.58) 56.40 (0.14) 50.58 (0.17)
PCL 13.52 (0.52) 64.79 (0.58) 5.35 (0.49) 64.70 (0.55) 60.47 (0.46) 51.56 (0.55) 56.08 (0.28) 50.99 (0.26)
ManyDG 14.24 (0.51) 65.98 (0.55) 6.06 (0.31) 64.66 (0.32) 62.38 (0.42) 53.19 (0.54) 57.81 (0.25) 52.34 (0.24)
SLDG 13.07 (0.50) 63.89 (0.60) 4.58 (0.44) 63.24 (0.60) 59.78 (0.49) 50.81 (0.53) 56.92 (0.14) 52.85 (0.16)
UdonCare 15.82 (0.33) 69.04 (0.42) 6.81 (0.27) 66.73 (0.48) 71.17 (0.35) 67.28 (0.39) 61.61 (0.10) 58.62 (0.25)
@@
Oracle 80.25 (0.12) 67.23 (0.31) 74.31 (0.25) 61.28 (0.22) 26.73 (0.12) 39.22 (0.18) 28.12 (0.11) 40.53 (0.16)
Base 68.54 (0.13) 47.65 (0.32) 66.94 (0.18) 53.13 (0.18) 21.51 (0.14) 30.83 (0.20) 20.07 (0.12) 31.52 (0.18)
DANN 75.32 (0.21) 60.82 (0.34) 69.63 (0.27) 53.43 (0.26) 21.84 (0.13) 34.51 (0.22) 24.05 (0.14) 35.21 (0.20)
CondAdv 76.81 (0.19) 64.18 (0.28) 71.48 (0.15) 55.62 (0.29) 22.81 (0.11) 36.48 (0.20) 26.13 (0.12) 37.35 (0.18)
MLDG 74.92 (0.22) 59.13 (0.31) 70.29 (0.27) 56.77 (0.16) 21.54 (0.15) 33.93 (0.21) 24.17 (0.13) 34.72 (0.19)
IRM 69.23 (0.19) 62.47 (0.33) 69.12 (0.14) 54.57 (0.18) 22.41 (0.14) 33.07 (0.22) 23.54 (0.15) 34.12 (0.21)
ManyDG 77.04 (0.20) 63.94 (0.30) 71.26 (0.19) 55.27 (0.19) 23.12 (0.12) 36.17 (0.21) 25.91 (0.13) 37.04 (0.20)
UdonCare 78.31 (0.18) 66.42 (0.32) 73.07 (0.33) 59.23 (0.14) 24.79 (0.10) 38.05 (0.19) 27.31 (0.11) 39.41 (0.17)
4.2 Main Results
Table 2 presents results on four classification tasks using MIMIC-III and MIMIC-IV. First, the performance gap between the Oracle and Base methods is substantial, showing the presence of considerable domain differences. Focusing on mortality prediction in MIMIC-III, DANN (Ganin et al., 2016) and MLDG (Li et al., 2018a), both relying on coarse domain partitions—show minimal improvements, likely due to difficulties in extracting consistent features from coarse partitions. PCL (Yao et al., 2022) exhibits a slight gain through proxy-to-sample relationships. Meanwhile, IRM (Arjovsky et al., 2019) and CondAdv (Isola et al., 2017) perform better by incorporating regularization or recurrent structures for binary temporal event prediction. Among clinical-specific baselines, ManyDG (Yang et al., 2023a) achieves the best results by leveraging mutual reconstruction, and SLDG (Wu et al., 2023) sees only modest improvement due to its reliance on the most recent admissions. Notably, UdonCare surpasses all baselines across all tasks. Specifically, UdonCare boosts the AUPRC score by around 5% for mortality in MIMIC-III, 8% for mortality in MIMIC-IV, and around 21% and 19% for readmission in MIMIC-III and MIMIC-IV, respectively. We further extend our experiments by adopting GAMENet (Shang et al., 2019b) and CGL (Lu et al., 2021) as the backbone (see Appendix E), which further strengthens the advantages of UdonCare.
4.3 More Quantitative Analysis
Effectiveness of Decomposition. Following the setting of Shen et al. (2022); Yang et al. (2023a), a linear classifier $d$ can be trained on the embedding to predict either (i) labels or (ii) domains. After training such a predictor, the cosine similarity can be calculated in terms of the learned weights to quantify the feature dimension overlaps. Note that $\mathbf{p}$, $\widetilde{\mathbf{r}}$, and $\mathbf{h}$ are normalized dimension-wise to ensure that each dimension is comparable. The results are shown in Table 3. In general, the third row has lower cosine similarities than the first two rows, which indicates that there is mostly non-overlap between feature dimensions predicting labels and domains. Moreover, the first two rows give relatively higher similarity and imply the domain and label information are separated from $\mathbf{p}$ into domain features $\tilde{\mathbf{r}}$ (scaling from $\mathbf{r}$) and invariant features $\mathbf{h}$. It provides the quantitative evidence that UdonCare stores domain and label information along distinct dimensions. Under the same setting, we also present the convergence process in detail (see Appendix G).
@@ 4.3 More Quantitative Analysis
Table 3: Cosine similarity of linear weights on $\mathbf{p}$, $\widetilde{\mathbf{r}}$, and $\mathbf{h}$
@@
* $W_{d}(\cdot)$ represents the learned linear weights. As an illustration, $W(\mathbf{p}\rightarrow\text{domains})$ denotes training a linear model on $\mathbf{p}$ to predict domain IDs, after which the weights are extracted. Cosine similarity scores are averaged across all classes and evaluated over 3 runs.

### v3

@@ 4.2 Mutual Forward Learning
Each domain can be viewed as the latent representation $\mathbf{r}$ sampled from a meta domain distribution $p(\cdot)$, so that we can identify $\mathbf{r}$ and then factorize $p(y|\mathbf{x})$ into $\int p(y|\mathbf{x},\mathbf{r})p(\mathbf{r}|\mathbf{x})\,d\mathbf{r}$ by approximating $q(\mathbf{r})\sim p(\mathbf{r}|\mathbf{x})$ given data samples $\mathbf{x}$. Subsequently, a domain encoder $p(\mathbf{r}|\mathbf{x})$ and a label predictor $p(y|\mathbf{x},\mathbf{r})$ are needed for inference. Here we parameterize the domain encoder $p(\mathbf{r}|\mathbf{x})$ as a network $g_{\theta}(\cdot)$ with parameter $\theta$. Since the pruned output matrix $\mathbf{M}$ (see Section 4.1) maps each training sample $\mathbf{x}$ to $\mathbf{m}$ (soft-label domain IDs), we apply $g_{\theta}(\cdot)$ to $\mathbf{m}$ to estimate the domain factor $\mathbf{r}:=g_{\theta}(\mathbf{m})$. Although $\mathbf{r}$ represents a probabilistic domain variable, we implement $g_{\theta}$ as a deterministic Multi-Layer Perceptron (MLP) for the prediction task. Next, we compute invariant features using a non-parametric function $h(\cdot):(\mathbf{r},\mathbf{p})\mapsto\mathbf{h}$, which fuses $\mathbf{r}$ (domain-level representation) and $\mathbf{p}$ (patient-level representation) as input features for the label predictor $p(y|\mathbf{x},\mathbf{r})$.
@@
$\mathcal{L}_{r}[g_{\theta}(\mathbf{m}),\bar{\mathbf{p}}]:=\mathrm{MSE}(\mathbf{r},\;\mathbb{E}[\mathbf{p}|\mathbf{m}])+\frac{\|\mathbf{r}_{\mu}-\mathbf{p}_{\mu}\|_{\mathcal{F}}^{2}}{\|\mathbf{p}_{\mu}\|_{\mathcal{F}}^{2}},$ (6)
@@
Invariant Feature Projection Learning. In Eq. (6), both $\mathbf{r}$ and $\mathbf{p}$ are rescaled into a shared vector space with comparable magnitudes. Hence, we can directly apply an orthogonal projection approach (as in early studies (Bousmalis et al., 2016; Shen et al., 2022; Yang et al., 2023a)) to obtain the invariant feature $\mathbf{h}$ by subtracting the parallel component of $\mathbf{p}$ in this shared vector space. We formalize this in $h(\cdot)$ as shown in Eq. (7):
@@ 4.3 Training and Inference
Iterative Training. To train UdonCare, we feed each data sample $\mathbf{x}$ into the hierarchy-pruning module to obtain its latent domain $\mathbf{m}$, and then perform two cross-reference steps under a mutual learning architecture. Rather than updating the model continuously in each epoch, we adopt an iterative training strategy, which prior studies (Cui et al., 2019; Sofiiuk et al., 2022) have shown can reduce training time while maintaining comparable predictive performance. [^1] For example, we set iterations $I=3$ and epochs $N=100$ by obtaining pretrained parameters for 40 epochs and then updating $\mathbf{M}$ iteratively every 20 epochs, yielding a total of 100 epochs. We iteratively update the model weights and regenerate domain assignments every 20 epochs in our experiment. Before each iteration, we reinitialize the parameters in $g_{\theta}(\cdot)$, because the input shape of $\mathbf{m}$ may change due to updated code-level embeddings. We also provide the pseudo-code of UdonCare in Appendix B.
Mutual Inference. After the orthogonal projection, we apply the network $q_{\xi}(\cdot)$ (operates on space $\mathbf{p}$) as a post-step to parameterize the label predictor $p(y|\mathbf{x},\mathbf{r})$; We can also parameterize $p(y|\mathbf{x})$ through the regular decoder network $d_{\eta}(\cdot)$ fed by patient embeddings from backbone $f_{\phi}(\cdot)$.
$p(y|\mathbf{x})\sim\hat{y}_{p}$ $=d_{\eta}(\mathbf{p})=d_{\eta}(f_{\phi}(\mathbf{x})),$
$p(y|\mathbf{x},\mathbf{r})\sim\hat{y}_{h}$ $=q_{\xi}(\mathbf{h})=q_{\xi}(h(g_{\theta}(\mathbf{m}),\mathbf{p})).$ (8)
@@ 5.1 Experiment Setups
Table 1: Performance comparison of four prediction tasks on MIMIC-III/MIMIC-IV. We report the average performance (%) and the standard deviation (in bracket) over 5 runs. We also bold for the best and underscore for the second best model.
@@
Oracle 16.95 (0.48) 70.61 (0.58) 8.46 (0.53) 68.92 (0.47) 74.42 (0.43) 69.74 (0.48) 69.37 (0.12) 68.47 (0.11)
Base 11.05 (0.65) 55.01 (0.98) 3.97 (0.47) 59.13 (0.76) 63.45 (0.84) 62.15 (0.73) 58.46 (0.24) 60.56 (0.31)
DANN 12.80 (0.52) 63.29 (0.69) 4.41 (0.45) 63.82 (0.46) 64.02 (0.61) 63.89 (0.75) 60.89 (0.11) 61.22 (0.23)
CondAdv 13.60 (0.52) 65.22 (0.53) 5.65 (0.62) 64.27 (0.68) 67.72 (0.41) 65.47 (0.50) 62.78 (0.14) 62.29 (0.19)
MLDG 13.25 (0.44) 64.07 (0.63) 4.75 (0.38) 62.75 (0.57) 66.00 (0.55) 65.59 (0.59) 60.20 (0.23) 62.24 (0.28)
IRM 13.50 (0.48) 65.41 (0.54) 4.14 (0.52) 62.36 (0.61) 67.07 (0.43) 65.93 (0.55) 61.81 (0.14) 61.84 (0.17)
PCL 13.70 (0.49) 64.55 (0.61) 5.35 (0.49) 64.70 (0.55) 65.03 (0.49) 66.36 (0.59) 61.03 (0.28) 63.04 (0.26)
ManyDG 14.10 (0.54) 66.27 (0.52) 6.06 (0.31) 64.66 (0.32) 68.63 (0.45) 67.03 (0.56) 63.48 (0.25) 64.07 (0.24)
SLDG 13.20 (0.47) 63.63 (0.56) 4.58 (0.44) 63.24 (0.60) 65.73 (0.53) 65.80 (0.50) 61.55 (0.14) 65.29 (0.16)
UdonCare 15.90 (0.36) 68.78 (0.40) 6.81 (0.27) 66.73 (0.48) 71.17 (0.32) 67.68 (0.36) 65.61 (0.10) 67.26 (0.25)
@@
Oracle 80.43 (0.14) 67.41 (0.29) 74.31 (0.25) 61.28 (0.22) 26.97 (0.13) 39.46 (0.27) 28.12 (0.11) 40.53 (0.16)
Base 68.83 (0.19) 57.31 (0.35) 66.94 (0.18) 53.13 (0.18) 21.32 (0.19) 31.08 (0.23) 20.07 (0.12) 31.52 (0.18)
DANN 75.18 (0.15) 60.55 (0.30) 69.63 (0.27) 53.43 (0.26) 22.08 (0.15) 34.27 (0.28) 24.05 (0.14) 35.21 (0.20)
CondAdv 76.59 (0.18) 64.42 (0.25) 71.48 (0.15) 55.62 (0.29) 23.09 (0.18) 36.27 (0.21) 26.13 (0.12) 37.35 (0.18)
MLDG 74.76 (0.17) 58.85 (0.33) 70.29 (0.27) 56.77 (0.16) 21.34 (0.17) 34.16 (0.26) 24.17 (0.13) 34.72 (0.19)
IRM 69.47 (0.16) 62.20 (0.30) 69.12 (0.14) 54.57 (0.18) 22.27 (0.16) 32.86 (0.23) 23.54 (0.15) 34.12 (0.21)
ManyDG 77.29 (0.14) 63.70 (0.28) 71.26 (0.19) 55.27 (0.19) 23.39 (0.14) 36.49 (0.27) 25.91 (0.13) 37.04 (0.20)
UdonCare 78.16 (0.18) 66.70 (0.29) 73.07 (0.33) 59.23 (0.14) 24.11 (0.18) 38.37 (0.26) 27.31 (0.11) 39.41 (0.17)
Predictive Tasks. We evaluate our approach on four representative tasks: (1) Mortality Prediction, which determines whether a patient will pass away by a specified time horizon after discharge. This is a binary classification task. (2) Readmission Prediction, which checks if a patient will be readmitted within a predefined window (e.g., next 15 days) following discharge. This is also framed as a binary classification. (3) Diagnosis Prediction, which forecasts the set of diagnoses (ICD-9-CM codes) for the patient’s next hospital visit based on prior visits. This requires multi-label classification. (4) Drug Recommendation, which suggests a set of medications (ATC-4 codes (Nahler and Nahler, 2009)) for the upcoming visit, also formulated as multi-label classification. These tasks reflect diverse clinical needs and provide a rigorous benchmark for evaluating DG methods.
Evaluation Metrics. Both readmission and mortality prediction are binary classification tasks, we calculate the Area Under the Precision-Recall Curve (AUPRC) and the Area Under the Receiver Operating Characteristic Curve (AUROC) scores due to the imbalanced label distribution. For drug recommendation, we evaluate predictions of all DG approaches by AUPRC and F1 scores, the same setting as ManyDG (Yang et al., 2023a) (i.e. $d<120$). For diagnosis prediction, we decide accurate prediction by weighted $\text{F}_{1}$ score as in Timeline and top-10 recall as in DoctorAI (Choi et al., 2016), since the former one measures the overall prediction on all classes (i.e. $d>4500$) and the latter one have concentration on positive code with low frequency.
@@
Table 2: Cosine similarity of linear weights on $\mathbf{p}$, $\widetilde{\mathbf{r}}$, and $\mathbf{h}$. Cosine similarity is averaged across classes and evaluated over 3 runs.
@@
* $W_{\text{d}}(\cdot)$ represents the learned linear weights. For example, $W_{\text{d}}(\mathbf{p}\rightarrow\text{domains})$ denotes training a linear model on $\mathbf{p}$ to predict domains.
Datasets & Data Split. We conduct experiments on two publicly available EHR databases, MIMIC-III and MIMIC-IV, which are widely used in clinical prediction (Johnson et al., 2016, 2023). MIMIC-III covers ICU admissions from 2001 to 2012, while MIMIC-IV spans 2008 to 2019. To avoid overlapping time ranges with MIMIC-III, we only retain patients from the years 2013201320192019 in MIMIC-IV. For each set of experiments, we extract 6,4976{,}497 and 49,55849{,}558 patients with multiple visits ($T\geq 2$) from both datasets as shown in Appendix E Table 5. Different from random data splitting, we evaluate our model under domain shifts induced by demographic and temporal factors. For the MIMIC-III dataset, patients with Medicare or Medicaid insurance are used as the source domain, while those with private insurance are treated as the target domain. Following the setting in SLDG (Wu et al., 2023) on the MIMIC-IV dataset, patients admitted after 2017 are assigned to the target domain, while all preceding patients are used as the source domain. We further divide the target subset into validation and test subsets using a fixed ratio of 20%:80%. Note that, to prevent information leakage, there is no patient overlap between the training and test sets.
Baselines. We compare UdonCare with two naïve baselines, five general DG baselines and two most recent clinical DG baselines: (1) Naïve Baselines: Oracle, trained backbone encoder directly on the target data, and Base, trained solely on the source data. The difference between metrics from Oracle (upper bound) and Base (lower bound) can show the distribution gaps between source and target data. (2) Typical DG Baselines: DANN (Ganin et al., 2016), CondAdv (Isola et al., 2017), MLDG (Li et al., 2018a), IRM (Arjovsky et al., 2019), and PCL (Yao et al., 2022). (3) Clinical DG Baselines: ManyDG (Yang et al., 2023a), and SLDG (Wu et al., 2023). Since drug recommendation and diagnosis prediction are multi-label classification, we drop PCL and SLDG in these two tasks due to their setting limitation. Note that, all baselines follow the same source/target definition as in the Data Split section, and the test set remains unseen to all models except Oracle. The results reported in following sections are the performance evaluated on the test set. More details of the baselines can be found in Appendix D, and UdonCare is implemented as described in Appendix F.
5.2 Main Results
Table 1 presents results on four classification tasks using MIMIC-III and MIMIC-IV. First, the performance gap between the Oracle and Base methods is substantial, showing the presence of considerable domain differences. DANN (Ganin et al., 2016) and MLDG (Li et al., 2018a), both relying on coarse domain partitions, show minimal improvements, likely due to difficulties in extracting consistent features from coarse partitions. PCL (Yao et al., 2022) exhibits a slight gain through proxy-to-sample relationships. Meanwhile, IRM (Arjovsky et al., 2019) and CondAdv (Isola et al., 2017) perform better by incorporating regularization or recurrent structures for predictions. Among clinical-specific baselines, ManyDG (Yang et al., 2023a) achieves the best results by leveraging mutual reconstruction, and SLDG (Wu et al., 2023) sees only modest improvement due to its reliance on the most recent admissions. Lastly, UdonCare surpasses over all baselines across all tasks. Specifically, UdonCare achieves a relative AUPRC improvement of around 11% (12%) for mortality prediction in MIMIC-III (MIMIC-IV), and 4% (3%) for readmission prediction on the MIMIC-III (MIMIC-IV) dataset.
Beyond the above results, we further extend our experiments by using GAMENet (Shang et al., 2019b) and CGL (Lu et al., 2021) as our backbone, and evaluate model generalization across hospitals on the eICU dataset (see Appendix G).
Figure 3: Domains among diseases of the circulatory system identified by UdonCare. The left table shows how the discovered domains correspond to specific ICD-9-CM diseases (codes) in the hierarchy, while the right figure is the scatter plot using t-SNE upon code-level embeddings.
5.3 Analysis of Discovered Domain
To illustrate how UdonCare behaves, we evaluate discovered domains on diagnosis prediction using MIMIC-III. Here, we focus on three kinds of heart related conditions from diseases of the circulatory system: heart failure (HF, 15 types), essential hypertension (EH. 3 types), and acute rheumatic fever (ARF, 7 types). In the left table of Figure 3, we can observe that the mid-level “heart failure”, high-level “hypertensive disease”, and several leaf-level nodes are chosen to represent these subtypes. These decisions reflect the same patterns of code-embedding geometry (right scatterplot plot): HF codes form a compact cluster, EH codes cluster even more tightly and lie near the HF group, while ARF codes are widely dispersed. Moreover, it is also consistent with medical knowledge: HF and EH are clinically cohesive, while ARF shows greater heterogeneity due to variable cardiac involvement. Thus, these results align with both data-driven and clinical logic.
5.4 More Quantitative Analyses
Effectiveness of Decomposition. Following the setting of (Shen et al., 2022; Yang et al., 2023a), a linear classifier $d$ can be trained on the embedding to predict either (i) labels or (ii) domains. After training such a predictor, the cosine similarity can be calculated in terms of the learned weights to quantify the feature dimension overlaps. Note that $\mathbf{p}$, $\widetilde{\mathbf{r}}$, and $\mathbf{h}$ are normalized dimension-wise to ensure that each dimension is comparable. The results are shown in Table 2. In general, the third row has lower cosine similarities than the first two rows, which indicates that there is mostly non-overlap between feature dimensions predicting labels and domains. Moreover, the first two rows give relatively higher similarity and imply the domain and label information are separated from $\mathbf{p}$ into domain features $\tilde{\mathbf{r}}$ (scaling from $\mathbf{r}$) and invariant features $\mathbf{h}$. It provides the quantitative evidence that UdonCare stores domain and label information along distinct dimensions. We provide more explanations in Appendix M. Under the same setting as Table 2, we also attach the convergence process in detail (see Appendix I).